Use the given vectors to compute u + v, u - v, and 5u - 2v. u = (9,-3), v = (1,7) U + V = U - V = 5u - 2y =
Answer #1

Given,
\\\mathrm{U = \left \langle 9,-3 \right \rangle = 9i -3j}\\ \mathrm{V = \left \langle 1,7 \right \rangle = 1i + 7j}\\ \\ \mathrm{1. \;\; U +V = 9i -3j + 1i + 7j ={\color{Blue} 10i + 4j = \left \langle 10,4 \right \rangle}}\\ \mathrm{2. \;\; U -V = 9i -3j - (1i + 7j) ={\color{Blue} 8i -10j = \left \langle 8,-10 \right \rangle}}\\ \mathrm{3. \;\; 5U -2V = 5(9i -3j) - 2(1i + 7j) =45i-15j-2i-14j={\color{Blue} 43i -29j = \left \langle 43,-29 \right \rangle}}\\

Please comment if you have any doubt in the solution. I will be happy to help you!